# Klear: Unified Multi-Task Audio-Video Joint Generation]Kling Team, Kuaishou Technology

### v1

@@ 5.2 Comparison with Existing Methods
Klear demonstrates robust audio–video joint generation. In Table 1, Klear achieves state-of-the-art performance, surpassing the two prior methods by a large margin. Cascaded approaches perform poorly due to error accumulation and strong dependence on upstream generation quality, while existing joint models exhibit only moderate audio–video consistency and noticeable unimodal degradation. In contrast, Klear attains substantially better A/V consistency and synchronization, which we attribute to the unified single-tower architecture and omni full attention mechanism.
Klear effectively guarantees unimodal performance. We then assess unimodal performance. As shown in Table 1, Klear delivers high audio and video quality in joint generation—surpassing cascaded and joint baselines by 34% and 18%, respectively—while multitask training yields more generalizable representations. Moreover, Klear outperforms specialized T2A and T2V models on their respective tasks, indicating that leveraging complementary audio–visual knowledge strengthens unimodal representations and further improves single-modality generation quality.
Klear consistently maintains a performance advantage across multiple tasks. Table 1 provides a comprehensive evaluation of Klear across a broad suite of tasks—including TI2AV, TI2V, T2V, and T2A—with comparisons to task-specialized baselines. Klear delivers consistently strong results, matching or surpassing specialized state-of-the-art models.
@@ 5.3 Qualitative Results
Lip-Sync Accuracy. Fig. 3 (a–b) highlights lip-sync performance. Klear achieves phoneme-level alignment between mouth movements and audio—covering mouth openings, lip–teeth shapes, and tongue positions, while Universe-1 and Ovi suffer from misalignment, delay, and clear audio–visual mismatch.
Emotional Expressiveness. From Fig. 3 (c), Klear generates characters with expressive emotions: facial cues (eyes, mouth curvature, muscle tension) are highly consistent with the audio’s affective tone (joy, sadness, excitement, lethargy), reflecting natural audio–visual fusion of prosody and dynamics. In contrast, Universe-1 and Ovi often produce distorted or emotional expressions.
@@
Singing and Rap Performance. As shown in Fig. 3, Klear yields performances where pitch, rhythm, and breath control are tightly aligned across audio and visuals—vibrato, melisma, and dynamic changes naturally match breathing patterns and facial expressions, consistent with human expectations of singing. In contrast, Universe-1 and Ovi show clear lip-sync failures, resulting in pronounced incongruity and reduced realism, especially for rap.
Audio–Visual Synchronization and Audio Overlapping. As shown in Fig. 3, Klear jointly generates background music and sound effects that are emotionally consistent with the video, with synchronized timing, realistic acoustics, and high fusion, thereby enhancing immersion. In contrast, baseline methods struggle to produce overlapping sounds and exhibit poor audio–visual consistency.
@@ 5.4 Ablation and Analysis
Architectural Effectiveness. To compare single- and dual-tower architectures for audio–video generation, we feed audio and video features into separate mm-DiT branches, with a randomly initialized audio tower due to the lack of a matching DiT backbone. Each block uses full attention and cross-attention for feature alignment. We also conduct an ablation with a pretrained audio tower. As shown in Table 2, the single-tower model outperforms the dual-tower variant in audio quality, video quality, and audio–video consistency, with visualizations confirming better cross-modal alignment. Although the pretrained audio tower converges faster, its performance is suboptimal due to a distribution mismatch with video features, hindering alignment.
@@ 6 Conclusion
We identify key failure modes in audio–video generation, such as asynchrony, lip–speech mismatch, and unimodal degradation, caused by suboptimal architectures, misaligned data, and single-task training. To address these, we propose Klear, a unified multi-task framework with a single-tower backbone and Omni-Full Attention for cross-modal interaction, alongside a progressive training strategy and an automated annotation pipeline. This pipeline produces a high-quality, annotated audio–video dataset for scalable training. Klear outperforms prior state-of-the-art on unimodal and joint audio-video benchmarks, becoming the superior model comparable to Veo 3. We hope this work could provide a clear direction and catalyze deeper research in audio–video generation.
7 Contributors
All contributors are listed as follows.
Jun Wang, Chunyu Qiang, Yuxin Guo, Yiran Wang, Feng Deng, Xijuan Zeng, Nan Li, Le Wang, Ming Wen, Zihan Li, Junjie Yan, Jingru Zhao, Zewen Song, Jiachen Zheng, Shiyao Wang, Xu Li, Zihao Ji, Zhongliang Liu, Xiaopeng Wang, Kang Yin, Teng Ma, Ziyu Zhang, Yuzhe Liang, Yushen Chen, Pengfei Cai, Haorui Zheng, Youjun Chen, Qianyue Hu, Jiahui Zhao, Yuejiao Wang, Liang Hou, Yuan Gao, Yan Zhou, Chen Zhang, Pengfei Wan.

### v2

@@ 5.2 Comparison with Existing Methods
Apollo demonstrates robust audio–video joint generation. In Table 1, Apollo achieves state-of-the-art performance, surpassing the two prior methods by a large margin. Cascaded approaches perform poorly due to error accumulation and strong dependence on upstream generation quality, while existing joint models exhibit only moderate audio–video consistency and noticeable unimodal degradation. In contrast, Apollo attains substantially better A/V consistency and synchronization, which we attribute to the unified single-tower architecture and omni-full attention mechanism.
Apollo effectively guarantees unimodal performance. We then assess unimodal performance. As shown in Table 1, Apollo delivers high audio and video quality in joint generation—surpassing cascaded and joint baselines by 34% and 18%, respectively—while multitask training yields more generalizable representations. Moreover, Apollo outperforms specialized T2A and T2V models on their respective tasks, indicating that leveraging complementary audio–visual knowledge strengthens unimodal representations and further improves single-modality generation quality.
Apollo consistently maintains a performance advantage across multiple tasks. Table 1 provides a comprehensive evaluation of Apollo across a broad suite of tasks—including TI2AV, TI2V, T2V, and T2A—with comparisons to task-specialized baselines. Apollo delivers consistently strong results, matching or surpassing specialized state-of-the-art models.
@@ 5.3 Qualitative Results
Lip-Sync Accuracy. Fig. 4 (a–b) highlights lip-sync performance. Apollo achieves phoneme-level alignment between mouth movements and audio—covering mouth openings, lip–teeth shapes, and tongue positions, while Universe-1 and Ovi suffer from misalignment, delay, and clear audio–visual mismatch.
Emotional Expressiveness. From Fig. 4 (c), Apollo generates characters with expressive emotions: facial cues (eyes, mouth curvature, muscle tension) are highly consistent with the audio’s affective tone (joy, sadness, excitement, lethargy), reflecting natural audio–visual fusion of prosody and dynamics. In contrast, Universe-1 and Ovi often produce distorted or emotional expressions.
@@
Singing and Rap Performance. As shown in Fig. 4, Apollo yields performances where pitch, rhythm, and breath control are tightly aligned across audio and visuals—vibrato, melisma, and dynamic changes naturally match breathing patterns and facial expressions, consistent with human expectations of singing. In contrast, Universe-1 and Ovi show clear lip-sync failures, resulting in pronounced incongruity and reduced realism, especially for rap.
Audio–Visual Synchronization and Audio Overlapping. As shown in Fig. 4, Apollo jointly generates background music and sound effects that are emotionally consistent with the video, with synchronized timing, realistic acoustics, and high fusion, thereby enhancing immersion. In contrast, baseline methods struggle to produce overlapping sounds and exhibit poor audio–visual consistency.
@@ 5.4 Ablation and Analysis
Architectural Effectiveness. To compare single- and dual-tower architectures for audio–video generation, we feed audio and video features into separate mm-DiT branches, with a randomly initialized audio tower due to the lack of a matching DiT backbone. Each block uses full attention and cross-attention for feature alignment. We also conduct an ablation with a pretrained audio tower. As shown in Table  2, the single-tower model outperforms the dual-tower variant in audio quality, video quality, and audio–video consistency, with visualizations confirming better cross-modal alignment. Although the pretrained audio tower converges faster, its performance is suboptimal due to a distribution mismatch with video features, hindering alignment.
@@ 6 Conclusion
We identify key failure modes in audio–video generation, such as asynchrony, lip–speech mismatch, and unimodal degradation, caused by suboptimal architectures, misaligned data, and single-task training. To address these, we propose Apollo, a unified multi-task framework with a single-tower backbone and Omni-Full Attention for cross-modal interaction, alongside a progressive training strategy and an automated annotation pipeline. This pipeline produces a high-quality, annotated audio–video dataset for scalable training. Apollo outperforms prior state-of-the-art on unimodal and joint audio-video benchmarks, becoming the superior model comparable to Veo 3. We hope this work could provide a clear direction and catalyze deeper research in audio–video generation.
Acknowledgement
We acknowledge the contributions from (sorted by first name): Boyuan Jiang, Chen Zhang, Feng Deng, Haorui Zheng, Jiachen Zheng, Jiahao Wang, Jiahui Zhao, Jingbin He, Jingke Li, Jingru Zhao, Junjie Yan, Kang Yin, Le Wang, Liang Hou, Lingyu Zou, Ming Wen, Nan Li, Peihan Li, Pengfei Cai, Pengfei Wan, Qianyue Hu, Shiyao Wang, Teng Ma, Xiaopeng Wang, Xiaoyu Shi, Xin Tao, Xu Li, Yan Zhou, Youjun Chen, Yu Zhao, Yuan Gao, Yuejiao Wang, Yun Li, Yushen Chen, Yuzhe Liang, Zewen Song, Zhongliang Liu, Zihan Li, Zihao Ji, Ziyang Yuan, and Ziyu Zhang.